OpenStudy (anonymous):

I'm not sure on these. I am to use rule for the square of a binomial. (5x^2-3)^2

OpenStudy (anonymous):

I may not get to be here for any help as I have to leave for school but I will check when I get back;) ty all in advance, Denise

OpenStudy (phi):

(5x^2-3)^2 is short for (5x^2-3)(5x^2-3) you multiply this out just like any binomial (a+b)(a+b) = a^2 + 2ab + a^2 If you know FOIL: 5x^2 * 5x^2 + 5x^2*(-3) + -3*5x^2 + -3*-3 5x^2 * 5x^2 = 5*5*x^2 * x^2= 25*x^4 = 25x^4 (multiply 5*5, add exponents for x^2*x^2) The whole thing simplifies to 25x^4 -30x^2+9